# SReferences

### v2

@@ B.2 Raw data collection and description
All six exchanges share similar data disclosure policies. Data are provided in an aggregated form over regular time intervals. For public users [^1] For institutional participants granted special access, the data disclosure frequency intensifies to 4 times per second., exchanges relay data to public channels every 500 milliseconds unless no market update has occurred, i.e., there has been neither a transaction nor an alteration to the order book in the last 500 millisecond; in this case, updates are not provided to the public channels. Table 2 provides the fields of data released by the exchanges. The “update time” field is the timestamp provided by the exchanges, which also functions as a unique identifier for a record. Fields “ask prices $1$ to $5$” and “bid prices $1$ to $5$” denote the topmost prices on the ask and bid sides, respectively, while the “ask volumes $1$ to $5$” and “bid volumes $1$ to $5$” denote the available quantities at the associated volume prices on the order book. The “volume” reflects the total trading volume of the derivative over the preceding 500 milliseconds, and the “turnover” measures the trading value in CNY (currency) for the same duration. “open interest” represents the aggregate count of unsettled derivative contracts, while “last price” indicates the most recent transaction price.
@@
In order to extract data from across China’s derivatives market, we implement an online data collection program that utilizes application programming interfaces (APIs) to connect with brokers. This program has been operational and amassing raw data since August 2022. The derivatives market in China operates on a membership basis. Individuals aiming to trade in the market must do so through brokers. These brokers, being members of the exchange, shoulder the responsibility of risk management for traders and clearing of trades. Exchanges engage solely with brokers and both market and trading data are conveyed through these brokers. For algorithmic traders, brokers offer APIs to facilitate data access. Figure 7 gives the number of weekly records accumulated through our data collection procedure. On average, daily records encompass 8 million entries for 700 futures instruments and 16 million entries for 10,000 option instruments. For futures, various maturities correspond to different instruments. In the case of options, combinations of differing maturities, strike prices, and types (call/put) lead to an array of instruments. Consequently, the instrument count for option data significantly surpasses that of futures data. We note that the options markets in China, while not as brisk as the futures market, demonstrates an upward trajectory in activity.
Figure 7: Weekly recorded numbers for futures and options.
@@ B.4 Potential research topics
[…]
@@ B.5 Standardized data
To create a standardized dataset conducive to accessibility, we have structured the data according to specific guidelines. These standards not only ensure uniformity across different instruments but also facilitate its ease of use by other researchers. The core attributes of our standardized dataset are as follows: (1) Standard Format: The dataset adheres to a consistent structure regardless of the instrument, making it seamless for other researchers to employ; (2) Market Representativeness: For this study, we constrain our focus to commodity future derivatives. The data spans a year, from 2022-08-01 to 2023-07-31, incorporating all products from the six aforementioned exchanges, except CFFEX. For each product, we select the most actively traded instrument, as gauged by its trading volume; (3) Trading Time Window: Notably, while some products undergo trading during the nighttime session, we’ve truncated the dataset to encapsulate only the daytime trading hours: 9:00 to 15:00; (4) Pricing Convention: We employ the mid-price as the standard price metric for each instrument. This mid-price is deduced by averaging the best bid and best ask prices. Subsequently, return series on 30-second basis are derived using the formula $R_{t}=\frac{P_{t+1}}{P_{t}}-1$. Returns that straddle across two trading days are flagged as missing values. In subsequent analyses, these can be supplanted by the mean return, which approximates to zero.
In summary, the dataset comprises $110,985$ observations across 55 distinct assets, with their classification detailed in Table 3. The classification methodology is based on the research report available at CITIC Securities Research Report. The dataset, along with the abbreviation code, is accessible via Google Drive, and higher-resolution data can be provided upon reasonable request to the authors.
@@ Appendix C Further simulation studies
We conduct additional simulation studies to investigate the performance of our testing procedures with a lower sample size and lower exceedance threshold. Specifically, we repeat the simulation design from Section 4 of the main paper but with a smaller sample size of $n=1,000$ and a lower exceedance threshold at the $0.9$-quantile. All other settings remain the same as in Section 4 of the main paper. The results are shown in Figure 8. The findings are consistent with those in Section 4 of the main paper, demonstrating that the permutation test with $\hat{\lambda}^{2}_{n}$ performs better than the $t$-test.
@@
Figure 8: Plot of p-values ($y$-axis) against empirical rank of test statistic $|\widehat{\lambda}_{n}^{(k)}(X,Y)|/K$ ($x$-axis); case $\lambda(X,Y)=0$ (Left two columns) and case $\lambda(X,Y)\neq 0$ (Right two columns); Red and blue points correspond to permutation tests using $\widehat{\lambda}_{n}^{1}$ and $\widehat{\lambda}^{2}_{n}$, respectively; purple points correspond to $t$-tests using $\widehat{\lambda}_{n}^{1}$. The underlying copula differs with the columns: the first and third columns are results for Gumbel copula and the second and fourth columns are results for Student’s $t$-copula. Rows correspond to different parameters for the copulas: for the Gumbel copula, $\theta\in\{1,1.2,1.5,2\}$ and for the Student’s $t$-copula, $\rho\in\{0.1,0.4,0.6,0.8\}$ with degrees of freedom $\nu$ fixed to be $4$.

### v3

@@ B.2 Raw data collection and description
All six exchanges share similar data disclosure policies. Data are provided in an aggregated form over regular time intervals. For public users, exchanges relay data to public channels every 500 milliseconds unless no market update has occurred, i.e., there has been neither a transaction nor an alteration to the order book in the last 500 millisecond; in this case, updates are not provided to the public channels. Table 3 provides the fields of data released by the exchanges. The “update time” field is the timestamp provided by the exchanges, which also functions as a unique identifier for a record. Fields “ask prices $1$ to $5$” and “bid prices $1$ to $5$” denote the topmost prices on the ask and bid sides, respectively, while the “ask volumes $1$ to $5$” and “bid volumes $1$ to $5$” denote the available quantities at the associated volume prices on the order book. The “volume” reflects the total trading volume of the derivative over the preceding 500 milliseconds, and the “turnover” measures the trading value in CNY (currency) for the same duration. “open interest” represents the aggregate count of unsettled derivative contracts, while “last price” indicates the most recent transaction price.
@@
In order to extract data from across China’s derivatives market, we implement an online data collection program that utilizes application programming interfaces (APIs) to connect with brokers. This program has been operational and amassing raw data since August 2022. The derivatives market in China operates on a membership basis. Individuals aiming to trade in the market must do so through brokers. These brokers, being members of the exchange, shoulder the responsibility of risk management for traders and clearing of trades. Exchanges engage solely with brokers and both market and trading data are conveyed through these brokers. For algorithmic traders, brokers offer APIs to facilitate data access. Figure 7 gives the number of monthly records accumulated through our data collection procedure. On average, daily records encompass 8 million entries for 700 futures instruments and 16 million entries for 10,000 option instruments. For futures, various maturities correspond to different instruments. In the case of options, combinations of differing maturities, strike prices, and types (call/put) lead to an array of instruments. Consequently, the instrument count for option data significantly surpasses that of futures data. We note that the options markets in China, while not as brisk as the futures market, demonstrates an upward trajectory in activity.
Figure 7: Monthly recorded numbers for futures and options.
@@ B.4 Potential research topics
With this dataset, a plethora of empirical research topics can be explored. Following a review of recent relevant literature on China’s derivatives market, we have summarized several potential research avenues: (1) Time-series dependence: Chen et al., (2021) investigated the interdependence between different sectors in China’s derivatives market using time-varying vector autoregressive (TV-VAR) and graphical models. Their study encompassed six sectors: petrochemicals, energy, softs, non-ferrous metals, oil and fats, and grains. Between 2004 and 2019, they identified that the energy and petrochemical sectors primarily drove spillover effects. Jia et al., (2016) delved into the temporal dependence between China’s agricultural futures markets and the US counterpart by examining lead-lag effects. Yang et al., 2021a analyzed the interplay between volatility in developed financial markets and the returns on China’s crude oil, by employing both conventional causality analysis and causality in quantiles; (2) Risk-contagions: Yang et al., 2021b investigated the risk spillover between the Chinese and global crude oil futures markets. By constructing upside and downside Value-at-Risk (VaR) connectedness networks, they observed a marked surge in risk spillover post the onset of the COVID-19 pandemic; (3) Market efficiency: Bohl et al., (2018) examined speculative activity against return volatility for crude oil products traded on DCE and CZCE. By defining speculative activity as the ratio of trading volume to open interest and executing Granger causality analysis, they discerned that heightened speculative activity tends to destabilize returns and amplify volatility. Conversely, Yang et al., (2020) assessed the efficiency disparity between the futures market of crude oil in the INE and the spot markets, predominantly contained within the Asia-Pacific region; (4) Market reaction to news: Li et al., 2022a explored the sentiment of relevant news and its repercussions on agricultural futures, by harnessing text analysis techniques. Additionally, Li et al., 2022b scrutinized the volatility spillover effects on crude oil products before and after the COVID-19 outbreak in 2022; (5) Trading strategy: Zhang et al., (2022) developed a trading bot based on the reinforcement learning algorithm to trade in the China’s derivatives market. This paper utilizes high-frequency data. However, the authors did not disclose the data source. Jiang et al., (2022) first studied the market-making strategy combined with reinforcement learning in China’s derivatives Market. Wu et al., (2022) study the momentum strategy in China’s derivatives market; (6) Market microstructure: So far, not many studies on the microstructure of the Chinese futures market can be found. However, in the meantime, this is an active topic in other markets such as cryptocurrency market (Almeida and Gonçalves, (2024)) US stock market (O’hara, (2015)). The high-frequency data we provide can be used to study the microstructure of China’s derivatives market; (7) Option pricing: Li et al., (2018) studied the correlation between past returns and the implied volatility spread (difference in implied volatility between call and put options with identical maturity and strike prices) for SSE 50 ETF options in China’s derivatives market. Contrary to the Black-Scholes Model’s prediction of a zero spread, they found a positive relationship with past returns, attributable to the momentum factor. Liu et al., (2023) studied the risk appetite and options-implied information.
@@ B.5 Standardized data
To create a standardized dataset conducive to accessibility, we have structured the data according to specific guidelines. These standards not only ensure uniformity across different instruments but also facilitate its ease of use by other researchers. The core attributes of our standardized dataset are as follows: (1) Standard Format: The dataset adheres to a consistent structure regardless of the instrument, making it seamless for other researchers to employ; (2) Market Representativeness: For this study, we constrain our focus to commodity future derivatives. The data spans a year, from 2022-08-01 to 2023-07-31, incorporating all products from the six aforementioned exchanges, except CFFEX. For each product, we select the most actively traded instrument, as gauged by its trading volume; (3) Trading Time Window: Notably, while some products undergo trading during the nighttime session, we’ve truncated the dataset to encapsulate only the daytime trading hours: 9:00 to 15:00; (4) Pricing Convention: We employ the mid-price as the standard price metric for each instrument. This mid-price is deduced by averaging the best bid and best ask prices. Subsequently, return series on a per-minute are derived using the formula $R_{t}=\frac{P_{t+1}}{P_{t}}-1$. Returns that straddle across two trading days are flagged as missing values. In subsequent analyses, these can be supplanted by the mean return, which approximates to zero.
In summary, the dataset comprises 6,615,735 observations across 55 distinct assets, with their classification detailed in Table 4. The classification methodology is based on the research report available at CITIC Securities Research Report. The dataset is publicly available on Zenodo at https://doi.org/10.5281/zenodo.15726389. Higher-resolution data can be provided upon reasonable request to the authors.
@@ Appendix C Further simulation studies
We conduct additional simulation studies to investigate the performance of our testing procedures with a lower sample size and lower exceedance threshold. Specifically, we repeat the simulation design from Section 4 of the main paper but with a smaller sample size of $n=1,000$ and a lower exceedance threshold at the $0.9$-quantile. All other settings remain the same as in Section 4 of the main paper. The results are shown in Figure 8. The findings are consistent with those in Section 4 of the main paper, demonstrating that the permutation test with $\hat{\lambda}^{2}_{n}$ performs better than the $t$-test. The power to reject the null hypothesis when data are generated under alternative hypothesis are summarized in Table 1 of the main paper. In addition, the empirical rejection rates when data are simulated under the null hypothesis are reported in Table 5 and Table 6. As expected, the rejection rates generally align with the nominal significance levels. However, for $n=1000$ and the $0.99$-quantile, only $1000\times 0.01=10$ effective samples are available, and all three proposed tests tend to exhibit biased rejection frequencies, especially the permutation-based test using the first estimator and the asymptotic test. Among them, the permutation-based test using the second estimator, $\hat{\lambda}^{2}_{n}$, yields rejection rates closest to the nominal level. This can be attributed to its data efficiency as explained in Section 4.2.
@@
Figure 8: Plot of p-values ($y$-axis) against empirical rank of test statistic $|\widehat{\lambda}_{n}^{(k)}(X,Y)|/K$ ($x$-axis); case $\lambda(X,Y)=0$ (Left two columns) and case $\lambda(X,Y)\neq 0$ (Right two columns); Circular (green) and triangular (yellow) points correspond to permutation tests using $\widehat{\lambda}_{n}^{1}$ and $\widehat{\lambda}^{2}_{n}$, respectively; Square (blue) points correspond to $t$-tests using $\widehat{\lambda}_{n}^{1}$. The underlying copula differs with the columns: the first and third columns are results for Gumbel copula and the second and fourth columns are results for Student’s $t$-copula. Rows correspond to different parameters for the copulas: for the Gumbel copula, $\theta\in\{1,1.5,2.5,3.0\}$ and for the Student’s $t$-copula, $\rho\in\{0.1,0.4,0.6,0.8\}$ with degrees of freedom $\nu$ fixed to be $4$.
Table 5: Empirical size (i.e., percentage of null hypotheses rejected when data are generated under the null hypothesis at a given significance level $\alpha^{*}$) of the three proposed statistical tests, at $\alpha^{*}=0.01$ and $\alpha^{*}=0.05$, across different sample sizes, $q$-quantile and Gumbel copula settings. The number of permutation is fixed as $P=1000$. For the Gumbel copula, $\theta\in\{1,1.5,2.5,3.0\}$. Entries with asterisk symbol denote significance levels that lie beyond three standard deviations from the mean.
$\theta$ $n$ $q$-quantile $\alpha^{*}=0.01$ $\alpha^{*}=0.05$
Perm. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{2}$ Asym. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{2}$ Asym. $\hat{\lambda}_{n}^{1}$
$1.1$ 1000 $q=0.90$ $1.0\pm 0.3$ $0.7\pm 0.3$ $0.8\pm 0.3$ $5.3\pm 0.7$ $4.1\pm 0.6$ $4.2\pm 0.6$
$q=0.99$ $0.2\pm 0.1^{*}$ $1.5\pm 0.4$ $0.0\pm 0.0^{*}$ $3.9\pm 0.6$ $5.4\pm 0.7$ $1.9\pm 0.4^{*}$
2500 $q=0.90$ $1.0\pm 0.3$ $0.6\pm 0.2$ $1.2\pm 0.3$ $5.6\pm 0.7$ $4.3\pm 0.6$ $5.8\pm 0.7$
$q=0.99$ $1.0\pm 0.3$ $1.0\pm 0.3$ $0.5\pm 0.2$ $5.4\pm 0.7$ $5.4\pm 0.7$ $5.1\pm 0.7$
5000 $q=0.90$ $0.7\pm 0.3$ $1.0\pm 0.3$ $0.8\pm 0.3$ $3.5\pm 0.6$ $4.9\pm 0.7$ $6.2\pm 0.8$
$q=0.99$ $0.6\pm 0.2$ $0.6\pm 0.2$ $0.6\pm 0.2$ $5.4\pm 0.7$ $4.5\pm 0.7$ $4.6\pm 0.7$
10000 $q=0.90$ $1.1\pm 0.3$ $1.1\pm 0.3$ $0.6\pm 0.2$ $5.0\pm 0.7$ $5.4\pm 0.7$ $4.6\pm 0.7$
$q=0.99$ $0.8\pm 0.3$ $0.9\pm 0.3$ $0.9\pm 0.3$ $5.4\pm 0.7$ $4.5\pm 0.7$ $5.3\pm 0.7$
$1.5$ 1000 $q=0.90$ $1.0\pm 0.3$ $1.5\pm 0.4$ $0.5\pm 0.2$ $5.0\pm 0.7$ $5.6\pm 0.7$ $3.9\pm 0.6$
$q=0.99$ $0.6\pm 0.2$ $1.2\pm 0.3$ $1.1\pm 0.3$ $3.2\pm 0.6^{*}$ $5.7\pm 0.7$ $6.3\pm 0.8$
2500 $q=0.90$ $1.0\pm 0.3$ $1.0\pm 0.3$ $0.5\pm 0.2$ $5.1\pm 0.7$ $4.7\pm 0.7$ $3.6\pm 0.6$
$q=0.99$ $0.9\pm 0.3$ $1.0\pm 0.3$ $0.7\pm 0.3$ $4.4\pm 0.6$ $4.6\pm 0.7$ $5.7\pm 0.7$
5000 $q=0.90$ $0.7\pm 0.3$ $0.9\pm 0.3$ $0.8\pm 0.3$ $4.6\pm 0.7$ $4.6\pm 0.7$ $5.0\pm 0.7$
$q=0.99$ $1.8\pm 0.4$ $1.3\pm 0.4$ $1.0\pm 0.3$ $5.6\pm 0.7$ $5.1\pm 0.7$ $5.8\pm 0.7$
10000 $q=0.90$ $1.8\pm 0.4$ $1.2\pm 0.3$ $0.9\pm 0.3$ $5.2\pm 0.7$ $4.7\pm 0.7$ $5.2\pm 0.7$
$q=0.99$ $1.3\pm 0.4$ $1.0\pm 0.3$ $1.3\pm 0.4$ $5.1\pm 0.7$ $4.9\pm 0.7$ $5.9\pm 0.7$
$2.0$ 1000 $q=0.90$ $0.8\pm 0.3$ $0.7\pm 0.3$ $0.5\pm 0.2$ $3.8\pm 0.6$ $3.8\pm 0.6$ $3.1\pm 0.5^{*}$
$q=0.99$ $0.1\pm 0.1^{*}$ $0.9\pm 0.3$ $1.5\pm 0.4$ $2.8\pm 0.5^{*}$ $4.3\pm 0.6$ $5.3\pm 0.7$
2500 $q=0.90$ $0.6\pm 0.2$ $0.9\pm 0.3$ $0.4\pm 0.2^{*}$ $4.7\pm 0.7$ $4.8\pm 0.7$ $3.2\pm 0.6^{*}$
$q=0.99$ $0.8\pm 0.3$ $0.8\pm 0.3$ $0.9\pm 0.3$ $4.1\pm 0.6$ $4.1\pm 0.6$ $4.7\pm 0.7$
5000 $q=0.90$ $0.7\pm 0.3$ $0.8\pm 0.3$ $0.5\pm 0.2$ $5.5\pm 0.7$ $5.4\pm 0.7$ $4.0\pm 0.6$
$q=0.99$ $0.8\pm 0.3$ $1.1\pm 0.3$ $0.6\pm 0.2$ $4.9\pm 0.7$ $4.7\pm 0.7$ $3.6\pm 0.6$
10000 $q=0.90$ $1.0\pm 0.3$ $0.9\pm 0.3$ $0.9\pm 0.3$ $4.7\pm 0.7$ $4.6\pm 0.7$ $4.9\pm 0.7$
$q=0.99$ $1.7\pm 0.4$ $0.9\pm 0.3$ $0.9\pm 0.3$ $5.2\pm 0.7$ $5.4\pm 0.7$ $5.2\pm 0.7$
$3.0$ 1000 $q=0.90$ $0.9\pm 0.3$ $0.4\pm 0.2^{*}$ $0.7\pm 0.3$ $4.5\pm 0.7$ $3.2\pm 0.6^{*}$ $4.4\pm 0.6$
$q=0.99$ $0.1\pm 0.1^{*}$ $1.6\pm 0.4$ $1.4\pm 0.4$ $3.2\pm 0.6^{*}$ $5.7\pm 0.7$ $4.9\pm 0.7$
2500 $q=0.90$ $0.6\pm 0.2$ $0.6\pm 0.2$ $0.5\pm 0.2$ $3.4\pm 0.6$ $4.1\pm 0.6$ $3.5\pm 0.6$
$q=0.99$ $0.8\pm 0.3$ $1.5\pm 0.4$ $1.1\pm 0.3$ $4.9\pm 0.7$ $4.4\pm 0.6$ $5.3\pm 0.7$
5000 $q=0.90$ $1.0\pm 0.3$ $0.4\pm 0.2^{*}$ $0.3\pm 0.2^{*}$ $3.3\pm 0.6^{*}$ $2.9\pm 0.5^{*}$ $3.7\pm 0.6$
$q=0.99$ $0.9\pm 0.3$ $1.8\pm 0.4$ $1.4\pm 0.4$ $4.9\pm 0.7$ $6.3\pm 0.8$ $5.3\pm 0.7$
10000 $q=0.90$ $0.6\pm 0.2$ $0.3\pm 0.2^{*}$ $1.0\pm 0.3$ $4.5\pm 0.7$ $3.8\pm 0.6$ $4.1\pm 0.6$
$q=0.99$ $0.8\pm 0.3$ $1.6\pm 0.4$ $1.4\pm 0.4$ $5.7\pm 0.7$ $5.2\pm 0.7$ $6.2\pm 0.8$
Table 6: Empirical size (i.e., percentage of null hypotheses rejected when data are generated under the null hypothesis at a given significance level $\alpha^{*}$) of the three proposed statistical tests, at $\alpha^{*}=0.01$ and $\alpha^{*}=0.05$, across different sample sizes, $q$-quantile and Student’s $t$-copula settings. The number of permutation is fixed as $P=1000$. For the Student’s $t$-copula, $\rho\in\{0.1,0.4,0.6,0.8\}$ with degrees of freedom fixed at $\nu=4$. Entries with asterisk symbol denote significance levels that lie beyond three standard deviations from the mean.
$\rho$ $n$ $q$-quantile $\alpha^{*}=0.01$ $\alpha^{*}=0.05$
Perm. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{2}$ Asym. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{2}$ Asym. $\hat{\lambda}_{n}^{1}$
$0.1$ 1000 $q=0.90$ $1.0\pm 0.3$ $1.0\pm 0.3$ $0.5\pm 0.2$ $5.1\pm 0.7$ $5.5\pm 0.7$ $4.6\pm 0.7$
$q=0.99$ $0.4\pm 0.2^{*}$ $1.2\pm 0.3$ $0.2\pm 0.1^{*}$ $2.9\pm 0.5^{*}$ $3.6\pm 0.6$ $2.8\pm 0.5^{*}$
2500 $q=0.90$ $1.4\pm 0.4$ $1.1\pm 0.3$ $0.6\pm 0.2$ $6.0\pm 0.8$ $5.7\pm 0.7$ $3.4\pm 0.6$
$q=0.99$ $1.1\pm 0.3$ $1.1\pm 0.3$ $1.3\pm 0.4$ $6.1\pm 0.8$ $4.3\pm 0.6$ $5.2\pm 0.7$
5000 $q=0.90$ $0.6\pm 0.2$ $1.3\pm 0.4$ $1.3\pm 0.4$ $4.1\pm 0.6$ $4.6\pm 0.7$ $4.6\pm 0.7$
$q=0.99$ $1.3\pm 0.4$ $0.9\pm 0.3$ $0.8\pm 0.3$ $5.8\pm 0.7$ $4.4\pm 0.6$ $4.9\pm 0.7$
10000 $q=0.90$ $0.7\pm 0.3$ $0.8\pm 0.3$ $1.3\pm 0.4$ $3.7\pm 0.6$ $5.0\pm 0.7$ $5.7\pm 0.7$
$q=0.99$ $1.2\pm 0.3$ $0.8\pm 0.3$ $0.8\pm 0.3$ $4.8\pm 0.7$ $5.0\pm 0.7$ $5.3\pm 0.7$
$0.4$ 1000 $q=0.90$ $1.0\pm 0.3$ $0.8\pm 0.3$ $1.3\pm 0.4$ $6.0\pm 0.8$ $5.6\pm 0.7$ $5.0\pm 0.7$
$q=0.99$ $0.0\pm 0.0^{*}$ $0.9\pm 0.3$ $0.7\pm 0.3$ $2.9\pm 0.5^{*}$ $4.1\pm 0.6$ $5.7\pm 0.7$
2500 $q=0.90$ $1.4\pm 0.4$ $0.8\pm 0.3$ $0.4\pm 0.2^{*}$ $4.4\pm 0.6$ $4.4\pm 0.6$ $4.8\pm 0.7$
$q=0.99$ $1.1\pm 0.3$ $1.1\pm 0.3$ $1.2\pm 0.3$ $4.7\pm 0.7$ $5.2\pm 0.7$ $5.3\pm 0.7$
5000 $q=0.90$ $1.4\pm 0.4$ $0.7\pm 0.3$ $0.8\pm 0.3$ $3.9\pm 0.6$ $4.2\pm 0.6$ $4.6\pm 0.7$
$q=0.99$ $0.8\pm 0.3$ $1.2\pm 0.3$ $1.4\pm 0.4$ $4.7\pm 0.7$ $5.6\pm 0.7$ $6.0\pm 0.8$
10000 $q=0.90$ $1.7\pm 0.4$ $0.9\pm 0.3$ $0.7\pm 0.3$ $5.7\pm 0.7$ $4.9\pm 0.7$ $4.4\pm 0.6$
$q=0.99$ $1.5\pm 0.4$ $0.8\pm 0.3$ $1.6\pm 0.4$ $5.4\pm 0.7$ $4.8\pm 0.7$ $4.1\pm 0.6$
$0.6$ 1000 $q=0.90$ $0.6\pm 0.2$ $1.2\pm 0.3$ $0.9\pm 0.3$ $5.0\pm 0.7$ $4.0\pm 0.6$ $3.7\pm 0.6$
$q=0.99$ $0.1\pm 0.1^{*}$ $0.9\pm 0.3$ $0.3\pm 0.2^{*}$ $3.7\pm 0.6$ $4.5\pm 0.7$ $3.7\pm 0.6$
2500 $q=0.90$ $1.3\pm 0.4$ $1.1\pm 0.3$ $0.5\pm 0.2$ $6.8\pm 0.8$ $5.0\pm 0.7$ $4.0\pm 0.6$
$q=0.99$ $1.5\pm 0.4$ $1.0\pm 0.3$ $0.9\pm 0.3$ $5.6\pm 0.7$ $4.6\pm 0.7$ $4.9\pm 0.7$
5000 $q=0.90$ $0.4\pm 0.2^{*}$ $0.1\pm 0.1^{*}$ $0.8\pm 0.3$ $4.3\pm 0.6$ $3.2\pm 0.6^{*}$ $5.0\pm 0.7$
$q=0.99$ $0.7\pm 0.3$ $1.2\pm 0.3$ $1.8\pm 0.4$ $5.1\pm 0.7$ $4.9\pm 0.7$ $5.5\pm 0.7$
10000 $q=0.90$ $0.8\pm 0.3$ $0.6\pm 0.2$ $0.7\pm 0.3$ $5.1\pm 0.7$ $3.6\pm 0.6$ $4.0\pm 0.6$
$q=0.99$ $1.2\pm 0.3$ $0.9\pm 0.3$ $1.1\pm 0.3$ $5.2\pm 0.7$ $5.6\pm 0.7$ $6.0\pm 0.8$
$0.8$ 1000 $q=0.90$ $1.0\pm 0.3$ $0.8\pm 0.3$ $1.2\pm 0.3$ $4.8\pm 0.7$ $4.2\pm 0.6$ $4.9\pm 0.7$
$q=0.99$ $0.2\pm 0.1^{*}$ $1.1\pm 0.3$ $1.8\pm 0.4$ $3.3\pm 0.6^{*}$ $5.0\pm 0.7$ $6.4\pm 0.8$
2500 $q=0.90$ $0.7\pm 0.3$ $0.4\pm 0.2^{*}$ $0.8\pm 0.3$ $4.4\pm 0.6$ $3.9\pm 0.6$ $4.4\pm 0.6$
$q=0.99$ $1.2\pm 0.3$ $0.6\pm 0.2$ $1.6\pm 0.4$ $5.3\pm 0.7$ $4.8\pm 0.7$ $7.3\pm 0.8$
5000 $q=0.90$ $0.6\pm 0.2$ $0.9\pm 0.3$ $0.6\pm 0.2$ $3.9\pm 0.6$ $4.3\pm 0.6$ $4.4\pm 0.6$
$q=0.99$ $1.1\pm 0.3$ $1.8\pm 0.4$ $0.8\pm 0.3$ $4.6\pm 0.7$ $5.6\pm 0.7$ $4.4\pm 0.6$
10000 $q=0.90$ $0.8\pm 0.3$ $0.8\pm 0.3$ $1.1\pm 0.3$ $3.7\pm 0.6$ $3.1\pm 0.5^{*}$ $4.7\pm 0.7$
$q=0.99$ $1.2\pm 0.3$ $0.9\pm 0.3$ $0.5\pm 0.2$ $5.3\pm 0.7$ $4.6\pm 0.7$ $4.8\pm 0.7$